Oldham Inc. conducts business in State M and State N, which both use the UDITPA three-factor formula to apportion income. State M’s corporate tax rate is 4.50 percent, and State N’s corporate tax rate is 7.00 percent. This year, Oldham had the following sales, payroll, and property (in thousands of dollars) in each state:

State M State N Total
Gross receipts from sales $ 3,600 $ 8,100 $ 11,700
Payroll expense 1,400 1,800 3,200
Property costs 1,500 1,600 3,100

If Oldham’s before-tax income was $3.6 million, compute its State M and State N tax. (Round all apportionment factors to 4 decimal places, e.g., 47.15% would be 0.4715. Enter "Apportionment percentage" and "Tax rate" in decimals and not in percentage. Enter your answers in dollars and not in millions or thousands of dollars.) STATE M STATE N

Total taxable income subject to apportionment:   

Apportionment Percentage:

State taxable income:

Tax Rate:

State Tax:
Answer #1

STATE MA STATE N Total taxable income Subject to apportionment sales factor (Note: 1) 0.3077 0.6923 payroll factor (Note: 2)Note: State M Slale N Total Gross Receipts from sales 11100 8100 3600 sales factor 0.3077 0.6923 ( 3600 ) 11700 (8100) 11 700Note: 3 Bled State M State N 3100 property costs 1500 1500 1600 1600 property factor 0.4839 9.5161 1600 Catoo)Note 4 Sum of factors Apportionment percentage Apportionment factors State M = 1.224 - 0.4097 State N = 1.7709 - 0.5903Note is Taxable income = 3600000 State taxable income - Total lexable x income Appos torment percentage of state Slate M = 36